KI and sucrose solution with 0.1M concentrations osmotic pressure of 0.465atm and 0.245atm.find the bant hadd KI and it's degree of dissociation.

Grade:12th pass

1 Answers

Vikas TU
4 years ago
Dear Student,
Please find below the solution to the asked query.
Osmotic pressure is given as:
π = iCRT
where π = osmotic pressure
i = van't Hoff factor
C = concentration
R = Gas constant
T = Temperature
So, for KI,
0.465 = i x 0.1 RT --------------------------------(1)
and for Sucrose,
0.245 = 1 x 0.1 RT -------------------------------(2)
(for sucrose, i = 1 as it doesn't dissociate into ions in solution)
Divide equation (1) by (2)
⇒0.4650/245=0.1iRT/0.1RT⇒
i = 1.897
For KI, degree of dissociation is:
α = (i−1)/(n−1)
or 
α = 1.897−1 / 2−1=0.897
or
α = 89.7%
So, van't Hoff factor = 1.897
Degree of dissociation = 89.7%
Hope this information clears your doubt about this topic. In case of any other doubt, do ask here on the portal and our experts will try to help you out as soon as possible.